if a=i-2j+k,b=4i+2J-k ,c=i+2j-k and a+kb is parallel to c then value of k is

Grade:10

2 Answers

siddhartha jain
11 years ago

a= i-2j+k, b=4i+2j-k

 

a+kb=(1+4k)i+(-2+2k)j+(1+k)k

since, a+kb is parallel to C,

hence the coefficients of i,j,k are same.

Hence on comparing, we get

 

K=2

Divyansh Garg
11 years ago

a + kb = i -2j + k +4ik + 2jk -k2 which can be written as-

(4k + 1)i + (2k - 2)j + (1-k)k

As a+k is parallel to c, the ratio of coefficients of i and j will be equal

:. (4k+1)/1 = (2k-2)/2

4k +1=k-1

k= -3/2

Hence solved